One of the most famous consequences of the relativity of time is the twin paradox.

Explain how the twin paradox comes about, and also explain why it does not break the

symmetry of special relativity


1
Expert's answer
2021-04-18T19:34:22-0400

Answer:

One of the twins embarks on a space journey to a distant star, moving at a speed close to the speed of light, and then returns. Due to the fact that time in the spacecraft-related frame of reference is slower than in the system associated with the Earth, his brother is much older than himself. The time in the ship's system ("\Delta"t ) is determined by the formula:

"{\displaystyle \Delta t=\Delta t_{0}{\sqrt {1-{\frac {v^{2}}{c^{2}}}}}}" ;

"\Delta t_0" - time on Earth,

"v-" is the speed of the ship,

"c-" is the speed of light.

For "v\approx c" the time in the ship system is small ("\Delta t\ll \Delta t_0" ).

On the other hand, he may believe that he himself remained inviolable, and the Earth first moved away from him and then returned. Then, in his view, it was his brother who made the journey almost at the speed of light, and therefore it was his brother who must remain much younger.


This connection can be explained as quantum entanglement. As Niels Bohr would explain, "any two objects that are created at the same time can affect each other." Therefore, the relationship between the twins can be called - quantum.

The first experiment to confirm this effect was set in 1971. Within the measurement error, the HAZ prediction was confirmed. Similar experiments (with increasing accuracy) also confirmed the correctness of the calculations (with an accuracy of 0.01%).